OpenStudy (anonymous):

Refer to the graph of y = sin x or y = cos x to find the exact values of x in the interval [0, 4pi] that satisfy the equation. 8 sin x = -8 I do not understand how to find the values on the graph? I know how to make the graph itself, but not how to do the specific problems on the graph.

OpenStudy (anonymous):

\[8\sin(x)=-8\] is the same as \[\sin(x)=-1\]

OpenStudy (anonymous):

there are only two values in the interval \[[0,4\pi]\] for which \[\sin(x0=-1\]

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=sin%28x%29+domain+0..4pi the minimum values of sine is \(-1\) and as you can see from the unit circle that \[\sin(x)=-1\] if \[x=\pi. 3\pi. 5\pi, ...\]

OpenStudy (anonymous):

Thank you so much, satellite73! The link to the graph was very helpful.

OpenStudy (anonymous):

yw